# Efficient and Scalable Graph Generation through Iterative Local Expansion

### v2

@@ 3.1 Graph Expansion
Given a graph $G=({\mathcal{V}},{\mathcal{E}})$ with $|{\mathcal{V}}|=n$ nodes and a cluster size vector ${\bm{v}}\in{\mathbb{N}}^{n}$ denoting the expansion size of each node, let $-~\tilde{G}(G,{\bm{v}})=(\tilde{{\mathcal{V}}},\tilde{{\mathcal{E}}})$ denote the expansion of $G$. It contains ${\bm{v}}[p]$ nodes, $v^{(p_{1})},\dots,v^{(p_{{\bm{v}}[p]})}$, for each node $v^{(p)}\in{\mathcal{V}}$ in the initial graph. As such, the expanded node set is given by $\tilde{{\mathcal{V}}}={\mathcal{V}}^{(1)}\cup\dots\cup{\mathcal{V}}^{(n)}$, where ${\mathcal{V}}^{(p)}=\{v^{(p_{i})}\mid 1\leq i\leq{\bm{v}}[p]\}$ for $1\leq p\leq n$. The edge set $-~\tilde{\mathcal{E}}$ includes all intracluster edges, $\{e^{\{p_{i},p_{j}\}}\mid 1\leq p\leq n,1\leq i<j\leq{\bm{v}}[p]\}$, as well as the cluster interconnecting edges, $\{e^{\{p_{i},q_{j}\}}\mid e^{\{p,q\}}\in{\mathcal{E}},v^{(p_{i})}\in{\mathcal{V}}^{(p)},v^{(q_{j})}\in{\mathcal{V}}^{(q)}\}$.
@@
Given a graph $-~\tilde{G}=(\tilde{{\mathcal{V}}},\tilde{{\mathcal{E}}})$ with $-~\tilde{m}=\left|\tilde{{\mathcal{E}}}\right|$ edges and an edge selection vector ${\bm{e}}\in\{0,1\}^{\tilde{m}}$, let $G(\tilde{G},{\bm{e}})=({\mathcal{V}},{\mathcal{E}})$ denote the refinement of $\tilde{G}$, with ${\mathcal{V}}=\tilde{{\mathcal{V}}}$ and $-~{\mathcal{E}}\subseteq\tilde{{\mathcal{E}}}$ such that the $i$-th edge $e^{(i)}\in{\mathcal{E}}$ if and only if ${\bm{e}}[i]=1$.
@@ Appendix B Perturbed Graph Expansion
Given a graph $G=({\mathcal{V}},{\mathcal{E}})$, a cluster size vector ${\bm{v}}\in{\mathbb{N}}^{n}$, a radius $r\in{\mathbb{N}}$, and a probability $0\leq p\leq 1$, the perturbed expansion $\tilde{G}$ is constructed as in Definition 1, and additionally for all distinct nodes $v^{(p)},v^{(q)}\in\tilde{{\mathcal{V}}}$ whose distance in $G$ is at most $r$, we add each edge $e^{\{p_{i},q_{j}\}}$ independently to $-~\tilde{{\mathcal{E}}}$ with probability $p$.
@@ E.1 Generative Modeling with Stochastic Differential Equations
$\operatorname*{\mathbb{E}}_{p({\bm{x}}_{0})}\operatorname*{\mathbb{E}}_{p({\bm{x}}_{t}\mid{\bm{x}}_{0})}\left[\left\|S_{{\bm{\theta}}}({\bm{x}}_{t})-\frac{{\bm{x}}_{t}-{\bm{x}}_{0}}{t^{2}}\right\|^{2}_{F}\right]$ (14)
@@
Instead of approximating the gradient directly, Karras et al. (2022) train a denoising model $D_{{\bm{\theta}}}({\bm{x}}_{t},t)$ to reconstruct ${\bm{x}}_{0}$ from ${\bm{x}}_{t}$. The joint training objective for all $t$ is given by:
@@
Here, $\lambda:[0,T]\rightarrow\mathbb{R}_{+}$ is a time-dependent weighting function, and $t$ is sampled such that $\ln(t)\sim N(P_{\text{mean}},P_{\text{std}})$ (details in Table 4). The gradient can then be recovered from the optimal denoising model $D_{{\bm{\theta}}}^{*}$ as $\nabla_{{\bm{x}}_{t}}\log p({\bm{x}}_{t})=D_{{\bm{\theta}}}^{*}({\bm{x}}_{t},t)$
@@ F.2 Local PPGN
${\color[rgb]{0.1484375,1,0.3203125}\definecolor[named]{pgfstrokecolor}{rgb}{0.1484375,1,0.3203125}\pgfsys@color@rgb@stroke{0.1484375}{1}{0.3203125}\pgfsys@color@rgb@fill{0.1484375}{1}{0.3203125}({\bm{h}}^{\prime})^{(1,1)}}=\gamma\left({\color[rgb]{0.1484375,1,0.3203125}\definecolor[named]{pgfstrokecolor}{rgb}{0.1484375,1,0.3203125}\pgfsys@color@rgb@stroke{0.1484375}{1}{0.3203125}\pgfsys@color@rgb@fill{0.1484375}{1}{0.3203125}{\bm{h}}^{(1,1)}},\phi\left({\color[rgb]{0.1484375,1,0.3203125}\definecolor[named]{pgfstrokecolor}{rgb}{0.1484375,1,0.3203125}\pgfsys@color@rgb@stroke{0.1484375}{1}{0.3203125}\pgfsys@color@rgb@fill{0.1484375}{1}{0.3203125}{\bm{h}}^{(1,1)}},{\color[rgb]{0.1484375,1,0.3203125}\definecolor[named]{pgfstrokecolor}{rgb}{0.1484375,1,0.3203125}\pgfsys@color@rgb@stroke{0.1484375}{1}{0.3203125}\pgfsys@color@rgb@fill{0.1484375}{1}{0.3203125}{\bm{h}}^{(1,1)}}\right)\oplus\left({\color[rgb]{1,0.8515625,0.3984375}\definecolor[named]{pgfstrokecolor}{rgb}{1,0.8515625,0.3984375}\pgfsys@color@rgb@stroke{1}{0.8515625}{0.3984375}\pgfsys@color@rgb@fill{1}{0.8515625}{0.3984375}{\bm{h}}^{(1,2)}},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\pgfsys@color@rgb@stroke{1}{0}{0}\pgfsys@color@rgb@fill{1}{0}{0}{\bm{h}}^{(2,1)}}\right)\oplus\left({\color[rgb]{0.40234375,0.671875,0.625}\definecolor[named]{pgfstrokecolor}{rgb}{0.40234375,0.671875,0.625}\pgfsys@color@rgb@stroke{0.40234375}{0.671875}{0.625}\pgfsys@color@rgb@fill{0.40234375}{0.671875}{0.625}{\bm{h}}^{(1,3)}},{\color[rgb]{0.80078125,0.6015625,1}\definecolor[named]{pgfstrokecolor}{rgb}{0.80078125,0.6015625,1}\pgfsys@color@rgb@stroke{0.80078125}{0.6015625}{1}\pgfsys@color@rgb@fill{0.80078125}{0.6015625}{1}{\bm{h}}^{(3,1)}}\right)\right)$
@@
${\color[rgb]{0.40234375,0.671875,0.625}\definecolor[named]{pgfstrokecolor}{rgb}{0.40234375,0.671875,0.625}\pgfsys@color@rgb@stroke{0.40234375}{0.671875}{0.625}\pgfsys@color@rgb@fill{0.40234375}{0.671875}{0.625}({\bm{h}}^{\prime})^{(1,3)}}=\gamma\left({\color[rgb]{0.40234375,0.671875,0.625}\definecolor[named]{pgfstrokecolor}{rgb}{0.40234375,0.671875,0.625}\pgfsys@color@rgb@stroke{0.40234375}{0.671875}{0.625}\pgfsys@color@rgb@fill{0.40234375}{0.671875}{0.625}{\bm{h}}^{(1,3)}},\phi\left({\color[rgb]{0.1484375,1,0.3203125}\definecolor[named]{pgfstrokecolor}{rgb}{0.1484375,1,0.3203125}\pgfsys@color@rgb@stroke{0.1484375}{1}{0.3203125}\pgfsys@color@rgb@fill{0.1484375}{1}{0.3203125}{\bm{h}}^{(1,1)}},{\color[rgb]{0.40234375,0.671875,0.625}\definecolor[named]{pgfstrokecolor}{rgb}{0.40234375,0.671875,0.625}\pgfsys@color@rgb@stroke{0.40234375}{0.671875}{0.625}\pgfsys@color@rgb@fill{0.40234375}{0.671875}{0.625}{\bm{h}}^{(1,3)}}\right)\oplus\left({\color[rgb]{1,0.8515625,0.3984375}\definecolor[named]{pgfstrokecolor}{rgb}{1,0.8515625,0.3984375}\pgfsys@color@rgb@stroke{1}{0.8515625}{0.3984375}\pgfsys@color@rgb@fill{1}{0.8515625}{0.3984375}{\bm{h}}^{(1,2)}},{\color[rgb]{0.3984375,0,0.80078125}\definecolor[named]{pgfstrokecolor}{rgb}{0.3984375,0,0.80078125}\pgfsys@color@rgb@stroke{0.3984375}{0}{0.80078125}\pgfsys@color@rgb@fill{0.3984375}{0}{0.80078125}{\bm{h}}^{(2,3)}}\right)\oplus\left({\color[rgb]{0.40234375,0.671875,0.625}\definecolor[named]{pgfstrokecolor}{rgb}{0.40234375,0.671875,0.625}\pgfsys@color@rgb@stroke{0.40234375}{0.671875}{0.625}\pgfsys@color@rgb@fill{0.40234375}{0.671875}{0.625}{\bm{h}}^{(1,3)}},{\color[rgb]{0.80078125,0.6015625,1}\definecolor[named]{pgfstrokecolor}{rgb}{0.80078125,0.6015625,1}\pgfsys@color@rgb@stroke{0.80078125}{0.6015625}{1}\pgfsys@color@rgb@fill{0.80078125}{0.6015625}{1}{\bm{h}}^{(3,1)}}\right)\right)$
@@
${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\pgfsys@color@rgb@stroke{0}{0}{1}\pgfsys@color@rgb@fill{0}{0}{1}({\bm{h}}^{\prime})^{(3,4)}}=\gamma\left({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\pgfsys@color@rgb@stroke{0}{0}{1}\pgfsys@color@rgb@fill{0}{0}{1}{\bm{h}}^{(3,4)}},\phi\left({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\pgfsys@color@rgb@stroke{0}{0}{1}\pgfsys@color@rgb@fill{0}{0}{1}{\bm{h}}^{(3,4)}},{\color[rgb]{0.3984375,0.69921875,1}\definecolor[named]{pgfstrokecolor}{rgb}{0.3984375,0.69921875,1}\pgfsys@color@rgb@stroke{0.3984375}{0.69921875}{1}\pgfsys@color@rgb@fill{0.3984375}{0.69921875}{1}{\bm{h}}^{(4,4)}}\right)\right)$

### v3

@@ 3.1 Graph Expansion
Given a graph $G=({\mathcal{V}},{\mathcal{E}})$ with $|{\mathcal{V}}|=n$ nodes and a cluster size vector ${\bm{v}}\in{\mathbb{N}}^{n}$ denoting the expansion size of each node, let $\tilde{G}(G,{\bm{v}})=(\tilde{{\mathcal{V}}},\tilde{{\mathcal{E}}})$ denote the expansion of $G$. It contains ${\bm{v}}[p]$ nodes, $v^{(p_{1})},\dots,v^{(p_{{\bm{v}}[p]})}$, for each node $v^{(p)}\in{\mathcal{V}}$ in the initial graph. As such, the expanded node set is given by $\tilde{{\mathcal{V}}}={\mathcal{V}}^{(1)}\cup\dots\cup{\mathcal{V}}^{(n)}$, where ${\mathcal{V}}^{(p)}=\{v^{(p_{i})}\mid 1\leq i\leq{\bm{v}}[p]\}$ for $1\leq p\leq n$. The edge set $\tilde{\mathcal{E}}$ includes all intracluster edges, $\{e^{\{p_{i},p_{j}\}}\mid 1\leq p\leq n,1\leq i<j\leq{\bm{v}}[p]\}$, as well as the cluster interconnecting edges, $\{e^{\{p_{i},q_{j}\}}\mid e^{\{p,q\}}\in{\mathcal{E}},v^{(p_{i})}\in{\mathcal{V}}^{(p)},v^{(q_{j})}\in{\mathcal{V}}^{(q)}\}$.
@@
Given a graph $\tilde{G}=(\tilde{{\mathcal{V}}},\tilde{{\mathcal{E}}})$ with $\tilde{m}=\left|\tilde{{\mathcal{E}}}\right|$ edges and an edge selection vector ${\bm{e}}\in\{0,1\}^{\tilde{m}}$, let $G(\tilde{G},{\bm{e}})=({\mathcal{V}},{\mathcal{E}})$ denote the refinement of $\tilde{G}$, with ${\mathcal{V}}=\tilde{{\mathcal{V}}}$ and ${\mathcal{E}}\subseteq\tilde{{\mathcal{E}}}$ such that the $i$-th edge $e^{(i)}\in{\mathcal{E}}$ if and only if ${\bm{e}}[i]=1$.
@@ Appendix B Perturbed Graph Expansion
Given a graph $G=({\mathcal{V}},{\mathcal{E}})$, a cluster size vector ${\bm{v}}\in{\mathbb{N}}^{n}$, a radius $r\in{\mathbb{N}}$, and a probability $0\leq p\leq 1$, the perturbed expansion $\tilde{G}$ is constructed as in Definition 1, and additionally for all distinct nodes $v^{(p)},v^{(q)}\in\tilde{{\mathcal{V}}}$ whose distance in $G$ is at most $r$, we add each edge $e^{\{p_{i},q_{j}\}}$ independently to $\tilde{{\mathcal{E}}}$ with probability $p$.
@@ E.1 Generative Modeling with Stochastic Differential Equations
$\operatorname*{\mathbb{E}}_{p({\bm{x}}_{0})}\operatorname*{\mathbb{E}}_{p({\bm{x}}_{t}\mid{\bm{x}}_{0})}\left[\left\|S_{{\bm{\theta}}}({\bm{x}}_{t})-\frac{{\bm{x}}_{0}-{\bm{x}}_{t}}{t^{2}}\right\|^{2}_{F}\right]$ (14)
@@
Instead of approximating the score directly, Karras et al. (2022) train a denoising model $D_{{\bm{\theta}}}({\bm{x}}_{t},t)$ to reconstruct ${\bm{x}}_{0}$ from ${\bm{x}}_{t}$. The joint training objective for all $t$ is given by:
@@
Here, $\lambda:[0,T]\rightarrow\mathbb{R}_{+}$ is a time-dependent weighting function, and $t$ is sampled such that $\ln(t)\sim N(P_{\text{mean}},P_{\text{std}})$ (details in Table 4). The score can then be recovered from the optimal denoising model $D_{{\bm{\theta}}}^{*}$ as $\nabla_{{\bm{x}}_{t}}\log p({\bm{x}}_{t})=(D_{{\bm{\theta}}}^{*}({\bm{x}}_{t},t)-{\bm{x}}_{t})/t^{2}$.
@@ F.2 Local PPGN
${\color[rgb]{0.1484375,1,0.3203125}\definecolor[named]{pgfstrokecolor}{rgb}{0.1484375,1,0.3203125}({\bm{h}}^{\prime})^{(1,1)}}=\gamma\left({\color[rgb]{0.1484375,1,0.3203125}\definecolor[named]{pgfstrokecolor}{rgb}{0.1484375,1,0.3203125}{\bm{h}}^{(1,1)}},\phi\left({\color[rgb]{0.1484375,1,0.3203125}\definecolor[named]{pgfstrokecolor}{rgb}{0.1484375,1,0.3203125}{\bm{h}}^{(1,1)}},{\color[rgb]{0.1484375,1,0.3203125}\definecolor[named]{pgfstrokecolor}{rgb}{0.1484375,1,0.3203125}{\bm{h}}^{(1,1)}}\right)\oplus\left({\color[rgb]{1,0.8515625,0.3984375}\definecolor[named]{pgfstrokecolor}{rgb}{1,0.8515625,0.3984375}{\bm{h}}^{(1,2)}},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}{\bm{h}}^{(2,1)}}\right)\oplus\left({\color[rgb]{0.40234375,0.671875,0.625}\definecolor[named]{pgfstrokecolor}{rgb}{0.40234375,0.671875,0.625}{\bm{h}}^{(1,3)}},{\color[rgb]{0.80078125,0.6015625,1}\definecolor[named]{pgfstrokecolor}{rgb}{0.80078125,0.6015625,1}{\bm{h}}^{(3,1)}}\right)\right)$
@@
${\color[rgb]{0.40234375,0.671875,0.625}\definecolor[named]{pgfstrokecolor}{rgb}{0.40234375,0.671875,0.625}({\bm{h}}^{\prime})^{(1,3)}}=\gamma\left({\color[rgb]{0.40234375,0.671875,0.625}\definecolor[named]{pgfstrokecolor}{rgb}{0.40234375,0.671875,0.625}{\bm{h}}^{(1,3)}},\phi\left({\color[rgb]{0.1484375,1,0.3203125}\definecolor[named]{pgfstrokecolor}{rgb}{0.1484375,1,0.3203125}{\bm{h}}^{(1,1)}},{\color[rgb]{0.40234375,0.671875,0.625}\definecolor[named]{pgfstrokecolor}{rgb}{0.40234375,0.671875,0.625}{\bm{h}}^{(1,3)}}\right)\oplus\left({\color[rgb]{1,0.8515625,0.3984375}\definecolor[named]{pgfstrokecolor}{rgb}{1,0.8515625,0.3984375}{\bm{h}}^{(1,2)}},{\color[rgb]{0.3984375,0,0.80078125}\definecolor[named]{pgfstrokecolor}{rgb}{0.3984375,0,0.80078125}{\bm{h}}^{(2,3)}}\right)\oplus\left({\color[rgb]{0.40234375,0.671875,0.625}\definecolor[named]{pgfstrokecolor}{rgb}{0.40234375,0.671875,0.625}{\bm{h}}^{(1,3)}},{\color[rgb]{0.80078125,0.6015625,1}\definecolor[named]{pgfstrokecolor}{rgb}{0.80078125,0.6015625,1}{\bm{h}}^{(3,1)}}\right)\right)$
@@
${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}({\bm{h}}^{\prime})^{(3,4)}}=\gamma\left({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}{\bm{h}}^{(3,4)}},\phi\left({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}{\bm{h}}^{(3,4)}},{\color[rgb]{0.3984375,0.69921875,1}\definecolor[named]{pgfstrokecolor}{rgb}{0.3984375,0.69921875,1}{\bm{h}}^{(4,4)}}\right)\right)$